# ViP2-CLIP: Visual-Perception Prompting with Unified Alignment for Zero-Shot Anomaly Detection

### v2

@@ Abstract
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
@@ 1 Introduction
Figure 1: Comparison between prior CLIP-based methods and ViP2-CLIP. ViP2-CLIP introduces ViP-Prompt to replace fixed class-name tokens with image-conditioned prompts that fuse global and local cues, and employs UTPA to perform a unified patch-level alignment.
Large‑scale vision–language models (VLMs) such as CLIP [23] have shown impressive zero‑shot recognition capabilities by pre‑training on vast image–text pairs. Inspired by this success, CLIP has been widely adopted for Zero‐Shot Anomaly Detection (ZSAD) [12], which seeks to detect anomalies automatically without any target domain training samples, offering a practical solution for data‑scarce scenarios. Unlike conventional image classification, which focuses on foreground semantics, anomaly detection emphasizes irregular regions on object surfaces. Consequently, existing CLIP‑based ZSAD methods typically construct two sets of text prompts—‘normal’ and ‘anomalous’, and align them with image features in a shared semantic space. As illustrated in Fig. 1 (a), text prompts are aligned with the global image feature to infer image-level anomaly score, while alignment with local patch features then generates pixel-level anomaly maps.
Currently, state-of-the-art ZSAD methods focus on prompt engineering to leverage CLIP’s generalization. Representative approaches such as WinCLIP [12], AnomalyCLIP [33], and VCP-CLIP [22] either craft semantically rich manual templates or learn prompt tokens to boost performance. However, manual templates demand extensive human effort and cannot guarantee comprehensive semantic coverage. Although some works [33] introduce object-agnostic learnable prompts to capture generic normal and abnormal concepts, they apply the same description across diverse anomaly types, hindering adaptability to complex semantic variations. Moreover, methods like CLIP-AD [6], AdaCLIP [4], FILo [8], and KANOCLIP [17] rely heavily on class names as prompt priors. Yet we find that CLIP’s segmentation quality remains acutely sensitive to semantically equivalent class-name variations, even after fine-tuning on auxiliary anomaly data (see Appendix A for detailed analysis). This likely stems from the limited scope of downstream adaptation, which cannot fully stabilize local alignment against label shifts. Such sensitivity constrains the class-name usage in prompts, highlighting the urgent need for robust, class-name agnostic but adaptive prompting strategies in ZSAD.
@@ 2 Related Work
ZSAD localizes abnormal regions without any target class samples during training. The advent of CLIP [23] has spurred notable advances in this task. WinCLIP [12] first ported CLIP to ZSAD with diverse handcrafted prompts. However, since CLIP is pretrained to align with general object semantics rather than abnormal patterns, its anomaly detection capability remains limited. To ease this burden, APRIL-GAN [5] and CLIP-AD [6] append shallow linear heads and fine-tune them on auxiliary datasets. AnomalyCLIP [33] removes handcrafted templates via learnable, object-agnostic prompts, boosting cross-domain generalization, while AdaCLIP [4] employs hybrid prompts to align text and image features better. VCP-CLIP [22] instead projects the image’s global feature into a fixed handcrafted prompt space, which enhances segmentation performance. By contrast, our ViP-Prompt fuses multi-scale visual cues with a learnable seed to generate dynamic, manual-free prompts, fully exploiting CLIP’s ZSAD performance, boosting both effectiveness and robustness in anomaly detection and localization.
@@ 3 Preliminary
CLIP consists of a text encoder $T(\cdot)$ and a visual encoder $F(\cdot)$, both implemented as mainstream multi-layer networks. Leveraging contrastive learning on large-scale image–text pairs, CLIP achieves superior zero-shot recognition. Given a class name $c$, we combine it with a text prompt template $G$ (e.g., ‘A photo of a [cls]’, where [cls] represents the category class-name). The resulting text is then fed into the text encoder $T(\cdot)$ to obtain the prompt embedding $g_{c}=T\bigl{(}G(c)\bigr{)}\in\mathbb{R}^{D}$. For an input image $x_{i}$, the visual encoder generates its global visual embedding $f_{i}\in\mathbb{R}^{D}$ and the local patch embeddings $f_{i}^{m}\in\mathbb{R}^{H\times W\times D}$. Specifically, given a category set $C$, CLIP computes the probability of image $x_{i}$ belonging to $c$ as follows:
@@
[…]
@@ 4.1 Visual-Perception Prompt (ViP-Prompt)
To condition prompts on the target object and thus adapt to complex detectionl scenarios, we map the global visual embedding $f$ extracted by the image encoder through a lightweight Meta-Net $h_{i}(\cdot)$ (Linear–ReLU–Linear) into the text embedding space, denoted as $\{a_{i}\}_{i=1}^{r}=h_{i}(f)$, where $r$ is the number of mapped tokens and $a_{i}\in\mathbb{R}^{C}$, then fuse it with static learnable vector to generate dynamic prompt token. Thus, we define the prompts’ structure as follows:
@@
where $z_{i}=o_{i}+a_{i}$. By projecting the image’s global feature into text embedding space, ICA adaptively generates normal and anomalous descriptors conditioned on the target object, eliminating the dependence on explicit class labels and enabling more robust cross‑modal alignment in privacy‑constrained settings.
@@
To further enhance the prompts’ fine-grained perceptual capacity, we introduce an attention-based interaction module between the text prompts and multi-scale local image features. Specifically, we project both text embeddings and local patch embeddings into a shared $C$-dimensional space. Let $G_{t}\in\mathbb{R}^{2\times C}$ be the prompt embeddings from the text encoder, and $F_{m}^{l}\in\mathbb{R}^{HW\times D}$ be the smoothed local visual embeddings from the $l$-th layer of the visual encoder, we then apply three learnable linear mappings to produce query $Q_{t}=G_{t}W_{q}$, key $K_{m}^{l}=F_{m}^{l}W_{k}$, and value $V_{m}^{l}=F_{m}^{l}W_{v}$, where $W_{q}\in\mathbb{R}^{C\times C}$, $W_{k},W_{v}\in\mathbb{R}^{D\times C}$, are learnable projection matrices.
@@
Figure 3: The visualization result of the attention maps from different prompts in our FGP module.
@@ 4.2 Unified Text-Patch Alignment (UTPA)
Task Datasets $|\mathcal{C}|$ CLIP WinCLIP APRIL-GAN AnomalyCLIP AdaCLIP ViP2-CLIP
OpenCLIP CVPR 2023 CVPRw 2023 ICLR 2024 ECCV 2024 -
Image-level (AUROC, AP, F1) MVTec AD 15 (66.5, 82.6, 86.7) (91.8, 96.5, 92.7) (86.1, 93.5, 90.4) (91.6, 96.4, 92.7) (90.1, 95.6, 92.3) (91.2, 96.0, 92.0)
VisA 12 (60.2, 66.2, 73.0) (78.1, 81.2, 78.2) (77.5, 80.9, 78.7) (82.0, 85.3, 80.4) (87.2, 89.7, 83.5) (88.5, 90.4, 84.8)
MPDD 6 (58.7, 69.7, 76.0) (61.5, 69.2, 77.5) (76.8, 83.0, 81.0) (77.5, 82.5, 80.4) (74.8, 78.6, 83.3) (79.7, 84.5, 82.4)
KSDD 1 (74.3, 55.3, 57.8) (92.4, 82.9, 77.7) (96.5, 91.2, 85.4) (97.8, 94.2, 89.7) (97.2, 92.6, 89.5) (98.1, 95.8, 93.2)
BTAD 3 (25.7, 49.8, 66.0) (68.2, 70.9, 67.8) (73.7, 69.7, 68.2) (88.2, 88.2, 83.8) (89.3, 96.5, 90.9) (95.0, 98.4, 94.7)
DAGM 10 (55.3, 43.2, 49.6) (91.8, 79.5, 75.7) (94.4, 83.9, 80.2) (97.7, 92.4, 90.1) (98.2, 92.3, 90.9) (98.5, 94.3, 92.6)
DTD-Synthetic 12 (52.3, 77.8, 85.1) (95.1, 97.7, 94.1) (85.6, 94.0, 89.1) (93.9, 97.2, 93.6) (96.3, 98.1, 95.5) (95.5, 98.1, 94.3)
AVERAGE - (56.1, 63.5, 70.6) (82.1, 82.1, 80.5) (84.4, 85.2, 81.9) (89.8, 90.9, 87.2) (90.4, 91.9, 89.4) (92.4, 93.9,90.6)
Pixel-level (AUROC, PRO, F1) MVTec AD 15 (35.6, 10.6, 6.9) (85.1, 64.6, 24.8) (87.6, 44.0, 43.3) (91.1, 81.4, 39.1) (89.6, 37.8, 45.1) (90.5, 87.1, 43.1)
VisA 12 (43.6, 14.0, 1.5) (79.6, 56.8, 9.0) (94.2, 86.6, 32.3) (95.5, 86.7, 28.3) (95.5, 56.8, 37.0) (95.4, 92.2, 33.6)
MPDD 6 (56.2, 27.3, 8.2) (71.2, 40.5, 15.4) (94.3, 83.8, 31.3) (96.5, 88.7, 34.2) (96.1, 60.3, 31.9) (97.2, 92.6, 35.9)
KSDD 1 (21.7, 2.6, 0.5) (92.8, 70.3, 15.8) (93.2, 84.1, 43.6) (98.1, 94.9, 56.5) (98.4, 53.0, 52.8) (98.5, 96.2, 53.0)
BTAD 3 (41.2, 10.8, 6.8) (72.7, 27.5, 18.5) (89.3, 68.7, 40.6) (94.2, 75.4, 49.7) (90.7, 22.3, 51.4) (95.6, 86.1, 52.7)
DAGM 10 (49.6, 12.4, 2.4) (87.6, 65.7, 12.7) (82.4, 66.0, 37.4) (95.6, 91.0, 58.9) (94.3, 42.5, 59.6) (97.5, 95.2, 61.2)
DTD-Synthetic 12 (43.8, 16.0, 3.7) (79.5, 51.4, 16.1) (95.2, 87.3, 67.4) (97.9, 92.0, 62.2) (98.5, 75.0, 71.8) (99.0, 96.5, 67.5)
AVERAGE - (41.7, 13.4, 4.3) (79.9, 52.6, 16.0) (90.9, 74.4, 42.2) (95.6, 87.2, 47.0) (94.7, 49.7, 49.9) (96.2,92.3, 49.6)
@@
Task Datasets $|\mathcal{C}|$ CLIP WinCLIP APRIL-GAN AnomalyCLIP AdaCLIP ViP2-CLIP
OpenCLIP CVPR 2023 CVPRw 2023 ICLR 2024 ECCV 2024 -
Image-level (AUROC, AP, F1) HeadCT 1 (67.8, 62.4, 70.9) (81.8, 80.2, 78.9) (89.1, 89.4, 82.1) (93.0, 91.1, 88.4) (94.0, 91.4, 90.1) (94.3, 93.9, 88.1)
BrainMRI 1 (72.2, 81.5, 76.5) (86.6, 91.5, 84.1) (89.4, 91.0, 88.2) (90.0, 92.1, 86.5) (94.3, 95.5, 92.2) (95.3, 96.7, 92.3)
Brain35H 1 (76.3, 77.7, 72.2) (79.9, 82.2, 74.0) (91.6, 92.1, 84.5) (93.4, 93.8, 86.4) (95.7, 95.8, 91.1) (95.8, 96.0, 90.1)
AVERAGE - (72.1, 73.9, 73.2) (82.8, 84.6, 79.0) (90.0, 90.8, 84.9) (92.1, 92.3, 87.1) (94.7, 94.2, 91.1) (95.1, 95.5, 90.2)
Pixel-level (AUROC, PRO, F1) ISIC 1 (43.2, 7.6, 44.0) (83.3, 55.1, 64.1) (89.4, 77.2, 71.4) (89.4, 78.4, 71.6) (91.3, 53.2, 75.5) (90.3, 82.3, 73.7)
CVC-ColonDB 1 (59.8, 30.4, 17.8) (64.8, 28.4, 21.0) (78.4, 64.6, 29.7) (81.9, 71.2, 37.5) (81.5, 64.3, 33.6) (82.5, 74.8, 36.2)
CVC-ClinicDB 1 (63.1, 33.8, 23.4) (70.3, 32.5, 27.2) (80.5, 60.7, 38.7) (82.9, 68.1, 42.4) (83.9, 65.7, 42.3) (86.3, 72.1, 44.9)
Endo 1 (59.7, 25.2, 28.9) (68.2, 28.3, 32.9) (81.9, 54.9, 44.8) (84.2, 63.4, 50.3) (86.0, 63.4, 51.8) (84.5, 63.7, 49.2)
Kvasir 1 (58.0, 18.3, 29.3) (69.7, 24.5, 35.9) (75.0, 36.3, 40.0) (79.0, 45.4, 46.2) (81.6, 49.1, 47.1) (81.9, 47.8, 47.3)
AVERAGE - (56.8, 23.1, 28.7) (71.3, 33.8, 36.2) (81.0, 58.7, 44.9) (83.5, 65.3, 49.6) (84.9, 59.1, 50.1) (85.1, 68.1, 50.3)

### v3

@@ 1 Introduction
Figure 1: Comparison between prior CLIP-based methods and ViP2-CLIP. ViP2-CLIP introduces ViP-Prompt to replace fixed class-name tokens with image-conditioned prompts that fuse global and local cues, and it first employs a unified patch-level alignment within training-based CLIP models.
Large‑scale vision-language models (VLMs) such as CLIP [26] have shown impressive zero‑shot recognition capabilities by pre‑training on vast image-text pairs. Inspired by this success, CLIP has been widely adopted for Zero‐Shot Anomaly Detection (ZSAD) [14], which seeks to detect anomalies automatically without any target domain training samples, offering a practical solution for data‑scarce scenarios. Unlike conventional image classification, which focuses on foreground semantics, anomaly detection emphasizes irregular regions on object surfaces. Consequently, existing CLIP‑based ZSAD methods typically construct two sets of text prompts–‘normal’ and ‘anomalous’, and align them with image features in a shared semantic space. As illustrated in Fig. 1 (a), text prompts are aligned with the global image feature to infer image-level anomaly score, while alignment with local patch features then generates pixel-level anomaly maps.
Currently, state-of-the-art ZSAD methods focus on prompt engineering to leverage CLIP’s generalization. Representative approaches such as WinCLIP [14], AnomalyCLIP [37], and VCP-CLIP [25] either craft semantically rich manual templates or learn prompt tokens to boost performance. However, manual templates demand extensive human effort and cannot guarantee comprehensive semantic coverage. Although some works [37] introduce object-agnostic learnable prompts to capture generic normal and abnormal concepts, they apply the same description across diverse anomaly types, hindering adaptability to complex semantic variations. Moreover, methods like CLIP-AD [6], AdaCLIP [4], FILo [10], KANOCLIP [19], and AA-CLIP [22] rely heavily on class names as prompt priors. Yet we find that CLIP’s segmentation quality remains acutely sensitive to semantically equivalent class-name variations, even after fine-tuning on auxiliary anomaly data (see Appendix A for detailed analysis). This likely stems from the limited scope of downstream adaptation, which cannot fully stabilize local alignment against label shifts. Such sensitivity constrains the class-name usage in prompts, highlighting the urgent need for robust, class-name agnostic but adaptive prompting strategies in ZSAD.
@@ 2 Related Work
ZSAD localizes abnormal regions without any target class samples during training. The advent of CLIP [26] has spurred notable advances in this task. WinCLIP [14] first ported CLIP to ZSAD with diverse handcrafted prompts. However, since CLIP is pretrained to align with general object semantics rather than abnormal patterns, its anomaly detection capability remains limited. To ease this burden, APRIL-GAN [5] and CLIP-AD [6] append shallow linear heads and fine-tune them on auxiliary datasets. AnomalyCLIP [37] removes handcrafted templates via learnable, object-agnostic prompts, boosting cross-domain generalization, while AdaCLIP [4] employs hybrid prompts to align text and image features better. VCP-CLIP [25] instead projects the image’s global feature into a fixed handcrafted prompt space, which enhances segmentation performance. AA-CLIP [22] adopted a two-stage training strategy that separately optimizes text and image encoders to encourage greater divergence in the text embedding space for improved anomaly discrimination. AF-CLIP [9] further integrated multi-scale feature aggregation to improve detection performance.
@@ 3 Preliminary
CLIP consists of a text encoder $T(\cdot)$ and a visual encoder $F(\cdot)$, both implemented as mainstream multi-layer networks. Leveraging contrastive learning on large-scale image-text pairs, CLIP achieves superior zero-shot recognition. Given a class name $c$, we combine it with a text prompt template $G$ (e.g., ‘A photo of a [cls]’, where [cls] represents the category class-name). The resulting text is then fed into the text encoder $T(\cdot)$ to obtain the prompt embedding $g_{c}=T\bigl(G(c)\bigr)\in\mathbb{R}^{D}$. For an input image $x_{i}$, the visual encoder generates its global visual embedding $f_{i}\in\mathbb{R}^{D}$ and the local patch embeddings $f_{i}^{m}\in\mathbb{R}^{H\times W\times D}$. Specifically, given a category set $C$, CLIP computes the probability of image $x_{i}$ belonging to $c$ as follows:
@@
where $\tau$ is a temperature hyperparameter and $\langle\cdot,\cdot\rangle$ denotes the cosine similarity.
@@ 4.1 Visual-Perception Prompt (ViP-Prompt)
To condition prompts on the target object and thus adapt to complex detection scenarios, we map the global visual embedding $f$ extracted by the image encoder through a lightweight Meta-Net $h_{i}(\cdot)$ (Linear–ReLU–Linear) into the text embedding space, denoted as $\{a_{i}\}_{i=1}^{r}=h_{i}(f)$, where $r$ is the number of mapped tokens and $a_{i}\in\mathbb{R}^{C}$, then fuse it with static learnable vector to generate dynamic prompt token. Thus, we define the prompts’ structure as follows:
@@
where $z_{i}=o_{i}+a_{i}$. By projecting the image’s global feature into the text embedding space, ICA adaptively generates normal and anomalous descriptors conditioned on the target object, eliminating the dependence on explicit class labels and enabling more robust cross‑modal alignment in privacy‑constrained settings.
@@
To further enhance the prompts’ fine-grained perceptual capacity, we introduce an attention-based interaction module between the text prompts and multi-scale visual features. Specifically, we project both text embeddings and local patch embeddings into a shared $C$-dimension space. Let $G_{t}\in\mathbb{R}^{2\times C}$ be the prompt embeddings from the text encoder, and $F_{m}^{l}\in\mathbb{R}^{HW\times D}$ be the smoothed local visual embeddings from the $l$-th layer of the visual encoder, we then apply three learnable linear mappings to produce query $Q_{t}=G_{t}W_{q}$, key $K_{m}^{l}=F_{m}^{l}W_{k}$, and value $V_{m}^{l}=F_{m}^{l}W_{v}$, where $W_{q}\in\mathbb{R}^{C\times C}$, $W_{k},W_{v}\in\mathbb{R}^{D\times C}$, are learnable projection matrices.
@@
Figure 3: Visualization results of the attention maps from different prompts in our FGP module.
@@ 4.2 Unified Text-Patch Alignment (UTPA)
Task Datasets $|\mathcal{C}|$ CLIP WinCLIP APRIL-GAN AnomalyCLIP AdaCLIP AA-CLIP ViP2-CLIP
OpenCLIP CVPR 2023 CVPRw 2023 ICLR 2024 ECCV 2024 CVPR 2025 -
Image-level
(AUROC, AP, F1)
MVTec AD 15 (66.5, 82.6, 86.7) (91.8, 96.5, 92.7) (86.1, 93.5, 90.4) (91.6, 96.4, 92.7) (90.1, 95.6, 92.3) (89.1, 94.7, 90.0) (91.2, 96.0, 92.0)
VisA 12 (60.2, 66.2, 73.0) (78.1, 81.2, 78.2) (77.5, 80.9, 78.7) (82.0, 85.3, 80.4) (87.2, 89.7, 83.5) (78.9, 82.3, 78.8) (88.5, 90.4, 84.8)
MPDD 6 (58.7, 69.7, 76.0) (61.5, 69.2, 77.5) (76.8, 83.0, 81.0) (77.5, 82.5, 80.4) (74.8, 78.6, 83.3) (56.9, 66.5, 74.5) (79.7, 84.5, 82.4)
KSDD 1 (74.3, 55.3, 57.8) (92.4, 82.9, 77.7) (96.5, 91.2, 85.4) (97.8, 94.2, 89.7) (97.2, 92.6, 89.5) (96.0, 89.2, 84.7) (98.1, 95.8, 93.2)
BTAD 3 (25.7, 49.8, 66.0) (68.2, 70.9, 67.8) (73.7, 69.7, 68.2) (88.2, 88.2, 83.8) (89.3, 96.5, 90.9) (93.4, 97.6, 93.9) (95.0, 98.4, 94.7)
DAGM 10 (55.3, 43.2, 49.6) (91.8, 79.5, 75.7) (94.4, 83.9, 80.2) (97.7, 92.4, 90.1) (98.2, 92.3, 90.9) (95.3, 87.5, 84.5) (98.5, 94.3, 92.6)
DTD-Synthetic 12 (52.3, 77.8, 85.1) (95.1, 97.7, 94.1) (85.6, 94.0, 89.1) (93.9, 97.2, 93.6) (96.3, 98.1, 95.5) (94.0, 98.0, 94.3) (95.5, 98.1, 94.3)
AVERAGE - (56.1, 63.5, 70.6) (82.1, 82.1, 80.5) (84.4, 85.2, 81.9) (89.8, 90.9, 87.2) (90.4, 91.9, 89.4) (86.7, 88.0, 86.5) (92.4, 93.9,90.6)
Pixel-level
(AUROC, PRO, F1)
MVTec AD 15 (35.6, 10.6, 6.9) (85.1, 64.6, 24.8) (87.6, 44.0, 43.3) (91.1, 81.4, 39.1) (89.6, 37.8, 45.1) (91.5, 86.5, 46.7) (90.5, 87.1, 43.1)
VisA 12 (43.6, 14.0, 1.5) (79.6, 56.8, 9.0) (94.2, 86.6, 32.3) (95.5, 86.7, 28.3) (95.5, 56.8, 37.0) (94.7, 82.7, 29.4) (95.4, 92.2, 33.6)
MPDD 6 (56.2, 27.3, 8.2) (71.2, 40.5, 15.4) (94.3, 83.8, 31.3) (96.5, 88.7, 34.2) (96.1, 60.3, 31.9) (96.0, 86.6, 26.5) (97.2, 92.6, 35.9)
KSDD 1 (21.7, 2.6, 0.5) (92.8, 70.3, 15.8) (93.2, 84.1, 43.6) (98.1, 94.9, 56.5) (98.4, 53.0, 52.8) (99.0, 92.2, 48.1) (98.5, 96.2, 53.0)
BTAD 3 (41.2, 10.8, 6.8) (72.7, 27.5, 18.5) (89.3, 68.7, 40.6) (94.2, 75.4, 49.7) (90.7, 22.3, 51.4) (94.2, 70.0, 45.1) (95.6, 86.1, 52.7)
DAGM 10 (49.6, 12.4, 2.4) (87.6, 65.7, 12.7) (82.4, 66.0, 37.4) (95.6, 91.0, 58.9) (94.3, 42.5, 59.6) (92.6, 79.7, 44.0) (97.5, 95.2, 61.2)
DTD-Synthetic 12 (43.8, 16.0, 3.7) (79.5, 51.4, 16.1) (95.2, 87.3, 67.4) (97.9, 92.0, 62.2) (98.5, 75.0, 71.8) (97.4, 88.9, 58.0) (99.0, 96.5, 67.5)
AVERAGE - (41.7, 13.4, 4.3) (79.9, 52.6, 16.0) (90.9, 74.4, 42.2) (95.6, 87.2, 47.0) (94.7, 49.7, 49.9) (95.0, 83.8, 42.5) (96.2,92.3, 49.6)
@@
Task Datasets $|\mathcal{C}|$ CLIP WinCLIP APRIL-GAN AnomalyCLIP AdaCLIP AA-CLIP ViP2-CLIP
OpenCLIP CVPR 2023 CVPRw 2023 ICLR 2024 ECCV 2024 CVPR 2025 -
Image-level
(AUROC, AP, F1)
HeadCT 1 (67.8, 62.4, 70.9) (81.8, 80.2, 78.9) (89.1, 89.4, 82.1) (93.0, 91.1, 88.4) (94.0, 91.4, 90.1) (88.1, 90.4, 80.6) (94.3, 93.9, 88.1)
BrainMRI 1 (72.2, 81.5, 76.5) (86.6, 91.5, 84.1) (89.4, 91.0, 88.2) (90.0, 92.1, 86.5) (94.3, 95.5, 92.2) (92.2, 94.5, 88.6) (95.3, 96.7, 92.3)
Brain35H 1 (76.3, 77.7, 72.2) (79.9, 82.2, 74.0) (91.6, 92.1, 84.5) (93.4, 93.8, 86.4) (95.7, 95.8, 91.1) (89.3, 90.6, 81.6) (95.8, 96.0, 90.1)
AVERAGE - (72.1, 73.9, 73.2) (82.8, 84.6, 79.0) (90.0, 90.8, 84.9) (92.1, 92.3, 87.1) (94.7, 94.2, 91.1) (89.9, 91.8, 83.6) (95.1, 95.5, 90.2)
Pixel-level
(AUROC, PRO, F1)
ISIC 1 (43.2, 7.6, 44.0) (83.3, 55.1, 64.1) (89.4, 77.2, 71.4) (89.4, 78.4, 71.6) (91.3, 53.2, 75.5) (91.6, 86.2, 78.1) (90.3, 82.3, 73.7)
CVC-ColonDB 1 (59.8, 30.4, 17.8) (64.8, 28.4, 21.0) (78.4, 64.6, 29.7) (81.9, 71.2, 37.5) (81.5, 64.3, 33.6) (80.6, 62.1, 32.3) (82.5, 74.8, 36.2)
CVC-ClinicDB 1 (63.1, 33.8, 23.4) (70.3, 32.5, 27.2) (80.5, 60.7, 38.7) (82.9, 68.1, 42.4) (83.9, 65.7, 42.3) (85.6, 66.3, 44.3) (86.3, 72.1, 44.9)
Endo 1 (59.7, 25.2, 28.9) (68.2, 28.3, 32.9) (81.9, 54.9, 44.8) (84.2, 63.4, 50.3) (86.0, 63.4, 51.8) (86.0, 66.5, 54.0) (84.5, 63.7, 49.2)
Kvasir 1 (58.0, 18.3, 29.3) (69.7, 24.5, 35.9) (75.0, 36.3, 40.0) (79.0, 45.4, 46.2) (81.6, 49.1, 47.1) (81.5, 48.4, 49.2) (81.9, 47.8, 47.3)
AVERAGE - (56.8, 23.1, 28.7) (71.3, 33.8, 36.2) (81.0, 58.7, 44.9) (83.5, 65.3, 49.6) (84.9, 59.1, 50.1) (85.1, 65.9, 51.6) (85.1, 68.1, 50.3)